# Back to the Color: Learning Depth to Specific Color Transformation for Unsupervised Depth Estimation

### v6

@@ Abstract
Virtual engines can generate dense depth maps for various synthetic scenes, making them invaluable for training depth estimation models. However, discrepancies between synthetic and real-world colors pose significant challenges for depth estimation in real-world scenes, especially in complex and uncertain environments encountered in unsupervised monocular depth estimation tasks. To address this issue, we propose Back2Color, a framework that predicts realistic colors from depth using a model trained on real-world data, thus transforming synthetic colors into their real-world counterparts. Additionally, we introduce the Syn-Real CutMix method for joint training with both real-world unsupervised and synthetic supervised depth samples, enhancing monocular depth estimation performance in real-world scenes. Furthermore, to mitigate the impact of non-rigid motions on depth estimation, we present an auto-learning uncertainty temporal-spatial fusion method (Auto-UTSF), which leverages the strengths of unsupervised learning in both temporal and spatial dimensions. We also designed VADepth, based on the Visual Attention Network, which offers lower computational complexity and higher accuracy than transformers. Our Back2Color framework achieves state-of-the-art performance on the Kitti dataset, as evidenced by improvements in performance metrics and the production of fine-grained details. This is particularly evident on more challenging datasets such as Cityscapes for unsupervised depth estimation. Code is available at https://github.com/BlueEg/back2color/.
Index Terms:
Back2Color, Syn-Real CutMix, auto-learning uncertainty, Visual Attention Depth Network
@@ I Introduction
In the real world, even after prolonged evolution, humans are still limited to qualitatively analyzing the depth of specific points in their field of view to assist in their actions[3]. Despite significant technological advances, the acquisition of precise and dense depth information remains a formidable challenge[4, 5]. Even with the refinement of advanced LiDAR systems[5], their data are inherently sparse, comprising only discrete depth points. Unsupervised depth estimation(UDE), powered by epipolar geometry[6] using purely visual approaches does not require additional annotation for depth information other than camera parameters. UDE eliminates the necessity of annotating depth labels for specific scenes. By utilizing binocular and motion disparity clues, it mimics the structure of human vision, making it efficient and cost-effective[7]. This is why many researchers and companies are actively pursuing this approach[8]. Monodepth2[9], the most renowned framework for UDE, has established itself as a classic and successful method. By reprojecting related frames(temporal and spatial) and regressing depth values. This adaptability enables Monodepth2 to obtain depth information for various scenes through adaptive fine-tuning, making it highly versatile. However, despite its strengths, Monodepth2 has several limitations. Firstly, the auto-masking for handling non-rigid moving objects, follows the rule: the pixel error after reprojection should be less than before reprojection. This pixel-level determination cannot actively detect motion[10] and cannot ensure effectiveness in smooth regions. Additionally, when applied to more complex datasets such as Cityscapes[11], it almost fails. Secondly, photometric reprojection error also fails in areas with no texture or with texture-less background objects, leading to depth prediction failure. Lastly, methods based on photometric reprojection matching lack prior knowledge of the objective structure. These issues make it challenging for UDE methods to be applied in real world.
@@
Figure 2: Illustration of our framework. Our Back2Color training follows a two-step optimization process within each iteration. In the first step, we use Optimizer 1 to train the Depth Net for supervised Syn-Real depth estimation and the Color Net to predict real colors from depths. Initially, ① we freeze the Depth Net and Color Net to infer both real and synthetic predicted depths, using synthetic depth data to predict Syn-to-Real Color. Next, ② we apply Syn-Real CutMix to generate new syn-real colors, and ③ optimize the Depth Net with synthetic ground truth and the Color Net on the real dataset. In the second step, with Optimizer 2, ④ we train for real-world unsupervised depth estimation, incorporating sky regularization loss and auto-learning uncertainty techniques. Note that the dotted line indicates inference only.
In this paper, our Back2Color learns the mapping process in the target domain (Real). This mapping is then applied to the source domain (Synthetic) to reduce the gaps between colors of the source and target domains. Additionally, we introduced an improved version of CutMix[20], termed Syn-Real CutMix, to effectively combine real-world, synthesized, and syn-to-real data in both batch and image spatial dimensions. By incorporating Syn-Real CutMix into our training process, we were able to enrich the training dataset with more diverse color-to-depth knowledge extracted from synthetic scenes. Furthermore, the joint training with diverse data sources enables the model to capture a wider range of color and depth variations present in real-world scenes, thus enhancing its robustness and generalization capability. Additionally, aiming to narrow the gap between predicted depths from real-world and synthesized depth, we addressed the issue of chaotic depth clouds (5) by introducing a Background Textureless Regularization Loss, also known as Sky Regularization Loss. Through joint training, our method enables the unsupervised monocular depth estimation to leverage knowledge from synthesized data, resulting in the generation of dense and sharp depths.
In pursuit of higher prediction accuracy and a more lightweight model compared to Transformer-based methods, we introduced the Visual Attention Network[21](VAN) as the backbone and equipped it with an efficient single-path high-resolution decoder, forming the new Visual Attention Depth Net(VADepth). Our design demonstrated more effective performance than Transformers. In our in-depth experiments, we found that Auto-masking is not always effective in identifying non-rigid moving objects with the same direction and speed. Additionally, it cannot handle vehicles moving in opposite directions, leading to significant errors in depth estimation. Therefore, we introduce the automatic learning uncertainty estimation fusion method[22], effectively combining the advantages of temporal-spatial unsupervised approaches, termed Auto-learning Uncertainty Temporal-Spatial Fusion(Auto-UTSF). This method thoroughly resolves the issue of non-rigid motion in unsupervised estimation, greatly improving the accuracy of depth estimation. To summarize, the key contributions of our work include:
  • To address the issue of color and texture discrepancies between synthetic and real-world datasets, we propose Back2Color. This framework stores the color information of specific scenes (target domain) within the model, enabling the reverse prediction of color images from depth data. Additionally, we introduce Syn-Real CutMix to harmonize the joint training of real-world and synthetic data, thereby enhancing the effectiveness of unsupervised depth estimation in real-world scenes.
  • We designed an efficient and lightweight Depth Net based on the Visual Attention Network. This network boasts fewer parameters and lower computational complexity compared to transformer-based networks, while delivering superior performance.
  • To more robustly recognize non-rigid moving objects and address occlusion issues in stereo systems, we enhance the method based on Auto-Uncertainty by introducing Auto-Spatial-Temporal Uncertainty Fusion, which leverages complementary spatial and temporal advantages in the unsupervised depth estimation process.
  • Experimental results demonstrate that our method surpasses all transformer-based models, achieving state-of-the-art (SOTA) metrics while maintaining manageable computational costs. Additionally, the visual boundaries of the estimated depth images are clearer.
@@
Figure 3: Color transformation for specific datasets: we showcase the capability of our Back2Color framework to effectively transform various synthetic datasets, including vKitti2[12], SHIFT[23], and UrbanSyn[24], into Syn2Real images. These images closely emulate the real-world styles of Kitti[25, 19] and Cityscapes[11], demonstrating high fidelity and credibility.
@@ II-A Unsupervised Depth Estimation
Unsupervised depth estimation aims to predict distances from stereo views or video sequences. Initial methods like those by Godard et al.[26] introduced disparity consistency loss, while Zhou et al.[27] pioneered depth and pose estimation simultaneously. Recent advancements[28, 29, 30, 31, 32, 33, 34, 35] refine network structures for higher accuracy. Techniques like PackNet[28] and CADepth[29] enhance geometric and appearance information propagation. HR-Depth[30] and DIFFNet[31] utilize HRNet for better resolution utilization. RA-Depth[32] employs HRNet for both encoding and decoding. In natural language processing, the Transformer architecture[36] has been successful, later adapted to computer vision tasks[37, 38]. It serves as the backbone for depth estimation[39, 40, 2, 1, 41], but faces challenges like computational demands and real-time constraints. Existing methods prioritize model architecture, often overlooking crucial obstacles and targets. Additionally, depth estimation models also face the issue of overfitting on specific datasets and a lack of continuous learning capability[42].
@@ II-C Style transfer in urban scenes
Research on style transfer for urban scene images has long been ongoing due to its importance in autonomous driving technology. Isola et al.[46] proposed a conditional GAN framework for generating urban scene images from semantic segmentation images of Cityscapes in 2017. However, generative models may produce random and plausible objects, disrupting the scene structure, as shown in their experiments. Atapour et al.[47] utilized style transfer and adversarial training to convert real-world color images into synthetic environment-style images. Huang et al.[48] applied GAN-based domain transfer for data augmentation in urban scene weather-time conversion. To minimize the gap between source and target domains, Gomez et al.[49] aligned colors across different domains in the Lab color space. Zheng et al.[50] used GANs to convert synthetic colors to real-world colors.
@@ III Methodology
In this section, we introduce the Back2Color framework, which masters the depth-to-color mapping within the target domain to train the Color Net (III-A1). Back2Color is enhanced by Real-Syn CutMix (III-A2) and a strategy for handling textureless background objects (III-A3). We then discuss the VADepth, based on the Visual Attention Network, designed for high performance and lightweight architecture (III-B). Lastly, we describe the Auto-learning Uncertainty Temporal-Spatial Fusion (Auto-UTSF) (III-C) for addressing non-rigid motion challenges. In this paper, Depth and Disparity are equivalent.
@@ III-A1 Transform depth to specific colors
Objectively speaking, predicting color from 3D structural information constitutes a one-to-many mapping, a non-standard conditional generative tasks, where the same 3D structure can correspond to a multitude of different color images. However, under specific conditions, this seemingly impossible task becomes feasible: Assuming that the color images in the target domain exhibit relatively high correlation. The colors in real word datasets like Cityscapes and Kitti are notably different, as shown in Fig.3. Therefore, in order to bridge the substantial gap between synthetic $\mathcal{S}\in${vKitti2[12](Vk),SHIFT[23](S), UrbanSyn[24]} and real-world $\mathcal{X}\in$Kitti[25, 19](K),Cityscapes[11](Cs)}datasets, we can learn the depth-to-color mapping in the target domain (real world color) and apply it to the source domain (synthetic color), thereby converting the source domain color to match those in the target domain. And then, we made the following attempts, train a Color Net $\mathcal{F}_{d2c}$ to predict color images $\tilde{\mathbf{I}}^{\mathcal{X}}$ from predicted depth images $\hat{\mathbf{D}}^{\mathcal{X}}$:
$\hat{\textbf{I}}^{\mathcal{X}}=\mathcal{F}_{d2c}(\hat{\textbf{D}}^{\mathcal{X}}|\boldsymbol{\theta}^{\mathcal{X}}),$ (1)
and utilizing the input of the Depth Net $\textbf{I}^{\mathcal{X}}$ for Supervision:
$\mathcal{L}^{d2c}=pe(\hat{\textbf{I}}^{\mathcal{X}},\textbf{I}^{\mathcal{X}}),$ (2)
Figure 5: Chaotic depth clouds: The green pixels can be matched across relevant frames based on the features of foreground obstacle, while the red pixels lacks effective features for matching, resulting in a cloud-like region in the depth image.
where $\boldsymbol{\theta}^{\mathcal{X}}$ represents the weights that store specific dataset color mapping information, such as Kitti and Cityscapes. Fortunately, we found that this prediction is feasible, neural networks are able to capture color information under specific target scenes. Next, we need to investigate whether we can employ $\boldsymbol{\theta}^{\mathcal{X}}$ to convert predicted $\hat{\textbf{D}}^{\mathcal{S}}$ in synthesized datasets into color that resemble the Kitti($\mathcal{X}$) color style: Using the specific Color Net to colorize predicted synthetic depth, as shown in Fig.3:
$\hat{\textbf{I}}^{\mathcal{S}\rightarrow\mathcal{X}}=\mathcal{F}_{d2c}(\hat{\textbf{D}}^{\mathcal{S}}|\boldsymbol{\theta}^{\mathcal{X}}).$ (3)
Figure 6: The proposed depth estimation module based on the Visual Attention Network. The encoder is composed of an ImageNet pretrained Visual Attention Network (VAN), generating multi-scale features. The VAN stages consist of the Large Kernel Attention (LKA) module, which is a pure CNN structure. This encoder balances performance and efficiency. The decoder features an HRNet structure, where each stage fuses multi-scales of features upward.
Our experiments have shown that by referencing color information from the target domain, we can learn its depth-to-color mapping and accurately predict color images in the source domain that closely resemble those in the target domain.
@@ III-A2 Syn-Real CutMix
CutMix[20] replaces removed pixels with patches from other images rather than simply removing them. This strategy eliminates uninformative pixels during training, enhancing efficiency, while maintaining the advantages of regional dropout to focus on non-discriminative obstacle parts. Additionally, introducing extra patches improves the model’s localization ability, requiring identification of obstacles from partial views. In our work, we draw inspiration from the CutMix method to achieve different objectives and effects. We aim to dynamically fuse Syn2Real Colors and Synthetic Colors within a single batch for supervised training, a technique we have named Syn-Real CutMix.
The goal of CutMix in our work is to generate a new training pair $(\tilde{\textbf{I}},\textbf{D})$. Unlike CutMix[20], our samples$(\textbf{I}^{\mathcal{S}},\hat{\textbf{I}}^{\mathcal{S}\rightarrow\mathcal{X}})$ are not two unrelated samples; instead, they correspond to different colored images of the same depth label $\textbf{D}^{\mathcal{S}}$. Therefore, we only merge color images without combining new labels. To increase the randomness of training samples. As shown in Fig.4, given image samples $(\textbf{I}^{\mathcal{S}},\hat{\textbf{I}}^{\mathcal{S}\rightarrow\mathcal{X}})$ as $(\textbf{I}_{1},\textbf{I}_{2})$, we firstly randomly swap $(\textbf{I}_{1},\textbf{I}_{2})$. Then, randomly crop a block from $\textbf{I}_{1}$ with a height and width greater than half of the original image, and replace it with the corresponding position in $\textbf{I}_{2}$:
$\tilde{\textbf{I}}=\textbf{M}\odot\textbf{I}_{1}+(1-\textbf{M})\odot\textbf{I}_{2},$ (4)
where $M_{i}\in\{0,1\}$ denotes a binary mask indicating where to drop out and fill in, and $\odot$ is element-wise multiplication.
Now, we combine Syn-Real enhanced images $\tilde{\textbf{I}}$ and Real world target image I into one batch for training. Then, we can train monocular depth estimation networks simultaneously on real-world, synthetic, and Syn-to-Real images within a batch. Spatial randomization also encourages the network to focus more on the image structure rather than its specific colors.
III-A3 Background textureless regularization loss
when analyzing and comparing unsupervised monocular depth and synthesized depth, we found not only significant gap in color images but also substantial gaps in the depth outputs generated by the model. In Fig.5, when the background of obstacle is textureless regions, these areas lack features for matching in sequences or stereo, making it challenging to accurately determine their depth. Consequently, these regions are prone to arbitrary depth assignments within the textureless area, leading to the formation of cloud-like regions in the depth map, we called them chaotic depth clouds. The main drawback of such cloud-like regions lies in their ability to significantly impact the overall distribution of depth values in the depth map, resulting in notable biases when calculating median variance during evaluation processes.
For chaotic depth clouds, they appear mostly within regions with the sky as background, such as trees, road signs, and buildings. Fortunately, the sky’s depth itself has a strong prior, that it is infinitely far away. Therefore, we predict the mask of the sky and constrain its value on the disparity map estimation by the network. This ensures that the network separates background and foreground obstacles:
$\mathcal{L}^{sky}=\frac{\sum|\textbf{M}_{sky}\times\hat{\textbf{D}}|}{\sum\textbf{M}_{sky}},$ (5)
where $\textbf{M}_{sky}$ is inferred from a pre-trained semantic segmentation network[52]. We also have an intuitive prior that the sky should only exist above in the view, so we added a decay coefficient $\textbf{G}\in\mathbb{R}^{W\times H}$ that decreases with the image height:
$\textbf{G}[..,j]=(H-j)/H,$ (6)
where $H$ is the height of $\textbf{D}_{t}$, so our final sky regularization loss function is:
$\mathcal{L}^{sky}=\frac{\sum\textbf{G}\times|\textbf{M}_{sky}\times\hat{\textbf{D}}|}{\sum\textbf{M}_{sky}}$ (7)
The proposed sky regularization loss effectively eliminates hazy depth while enhancing the recovery of delicate objects.
@@ III-A3 Background textureless regularization loss
III-B Visual Attention Depth Net
Every high-performing depth estimation network requires an effective and robust backbone to extract hierarchical features from natural images. Commonly used backbones include ResNet[53] (Monodepth2[9], HRDepth[30], CADepth[29]), HRNet[54] (DIFFNet[31], RA-Depth[32]), and Transformer-based backbones like ViT[37], MPViT[41](TSUDepth[22], MonoViT[2]). With the evolution from ResNet to HRNet and then to Transformer-based approaches, the performance of depth estimation continues to improve, but accompanied by increasing computational complexity and params. We consistently believe that depth estimation tasks, especially those based on street scenes, need to consider the requirements of their usage scenarios (real-time, lightweight). Hence, we aim to find a more efficient and effective backbone than Transformer methods.
Research on balancing the performance and efficiency of CNN and Transformer models has been ongoing, and we have discovered the VAN[21] (Visual-Attention-Network) model. The VAN[21] model employs large kernel convolutions instead of self-attention mechanisms to capture long-range dependencies. As shown in Fig.6, It decomposes the large kernel convolution into a spatial local convolution (depth-wise convolution), a spatial long-range convolution (depth-wise dilation convolution), and a channel convolution (1×1 convolution) to reduce computational complexity. This design not only avoids treating 2D images as 1D sequences but also achieves stronger accuracy on ImageNet while reducing computational complexity and params.
Due to its outstanding performance and efficiency, we have introduced the VAN model into unsupervised monocular depth estimation tasks, and adapt to some fine-tuning. The most significant change is that, to reduce computational complexity, VAN downsamples the input resolution to $2^{-2}$ during the first stem stage. While this downsampling is acceptable for classification tasks, it poses challenges for high-resolution recovery in image regression tasks. Therefore, without altering the original VAN structure, we have added an additional stem head to acquire features at $2^{-1}$ resolution, as shown in Fig.6. After extracting features from various VAN stages, we equipped them with a high-resolution depth decoder[32] to output the final depth. During implementation, we observed no difference between outputting a single-size depth image and outputting multi-scale depth images. Therefore, our network only outputs depth maps with the same size as the input.
Figure 8: Temporal reprojection involves reprojecting the previous and subsequent frames to the target frame, while spatial reprojection involves reprojecting the left or right frame to the target frame.

### v7

# Back2Color: Domain-Adaptive Synthetic-to-Real Monocular Depth Estimation for Dynamic Traffic Scenes

@@ Abstract
Accurate monocular depth estimation is a fundamental component of vision-based perception systems in intelligent transportation applications. Despite recent progress, unsupervised monocular approaches still suffer from significant performance degradation in real-world traffic scenes due to synthetic-to-real domain gaps and the presence of dynamic, non-rigid objects such as vehicles and pedestrians. In this paper, we propose Back2Color, a robust unsupervised monocular depth estimation framework that addresses these challenges through domain adaptation and uncertainty-aware fusion. Specifically, Back2Color proposes a bidirectional depth-to-color transformation strategy that learns appearance mappings from real-world driving data and applies them to synthetic depth maps, thereby constructing training samples with realistic color appearance and paired synthetic depth. In this way, the proposed approach effectively reduces the domain gap between simulated and real traffic scenes, enabling the depth prediction network to learn more stable and generalizable priors. To further improve robustness under dynamic environments, we propose an auto-learning uncertainty temporal-spatial fusion (Auto-UTSF) module, which adaptively fuses complementary temporal and spatial cues by estimating pixel-wise uncertainty, enabling reliable depth prediction in the presence of moving objects and occlusions. Extensive experiments on challenging urban driving benchmarks, including KITTI and Cityscapes, demonstrate that the proposed method consistently outperforms existing unsupervised monocular depth estimation approaches, particularly in dynamic traffic scenarios, while maintaining high computational efficiency.
@@ I Introduction
In the real world, even after prolonged evolution, humans are still limited to qualitatively analyzing the depth of specific points in their field of view to assist in their actions[23]. Despite significant technological advances, the acquisition of precise and dense depth information remains a formidable challenge[15, 36, 30, 50, 39, 37]. Even with the refinement of advanced LiDAR systems[36], their data are inherently sparse, comprising only discrete depth points. Unsupervised depth estimation(UDE), powered by triangulation[42] using purely visual approaches does not require additional annotation for depth information other than camera parameters. UDE eliminates the necessity of annotating depth labels for specific scenes. By utilizing binocular and motion disparity clues, it mimics the structure of human vision, making it efficient and cost-effective[3]. This is why many researchers and companies are actively pursuing this approach[28]. Monodepth2[14], the most renowned framework for UDE, has established itself as a classic and successful method. By reprojecting related frames(temporal and spatial) and regressing depth values. This adaptability enables Monodepth2 to obtain depth information for various scenes through adaptive fine-tuning, making it highly versatile. However, despite its strengths, Monodepth2 has several limitations. Firstly, the auto-masking for handling non-rigid moving objects, follows the rule: the pixel error after reprojection should be less than before reprojection. This pixel-level determination cannot actively detect motion[24] and cannot ensure effectiveness in smooth regions. Additionally, when applied to more complex datasets such as Cityscapes[7], it almost fails. Secondly, photometric reprojection error also fails in areas with no texture or with texture-less background objects, leading to depth prediction failure. Lastly, methods based on photometric reprojection matching lack prior knowledge of the objective structure. These issues make it challenging for UDE methods to be applied in real world.
@@
Figure 2: Illustration of the Back2Color training strategy. In each iteration, the first stage jointly optimizes the depth estimation network and the depth-to-color network using real-world data with unsupervised photometric and color restoration losses. In the second stage, the depth-to-color network is fixed and applied to transform synthetic depth maps into realistic color images, enabling the depth estimation network to be trained with virtual depth supervision under a reduced domain discrepancy.
In this paper, we propose the Back2Color framework to address the appearance discrepancy between synthetic and real-world data in unsupervised monocular depth estimation. Specifically, Back2Color introduces a bidirectional depth-to-color transformation strategy that learns appearance mappings from real-world driving scenes and applies them to synthetic depth maps. In this way, Back2Color constructs training pairs that combine realistic color appearance with corresponding synthetic depth, effectively narrowing the domain gap between simulated and real traffic scenes.
To further enable effective joint training across heterogeneous data sources, we introduce an enhanced CutMix strategy, termed Syn-Real CutMix [59], which mixes synthetic and syn-to-real images at both batch and spatial levels. This strategy enriches the training data with diverse color–depth associations derived from synthetic scenes while preserving real-world appearance characteristics. By jointly training on these complementary data sources, the proposed framework allows the depth estimation network to learn more stable and generalizable depth prediction priors, leading to improved robustness in real traffic environments.
In addition, we observe that existing auto-masking mechanisms often fail to reliably identify non-rigid moving objects in traffic scenes, especially when vehicles move at similar speeds or in opposite directions, which leads to erroneous depth predictions. To address this issue, we introduce an automatic uncertainty learning fusion strategy that combines temporal and spatial cues, termed Auto-learning Uncertainty Temporal-Spatial Fusion (Auto-UTSF) [69]. By explicitly modeling uncertainty, Auto-UTSF adaptively integrates complementary information from stereo and temporal supervision, enabling more reliable depth estimation under dynamic and occluded conditions.
To summarize, the main contributions of this work are as follows:
  • We propose Back2Color, a synthetic-to-real adaptation framework that learns scene-specific color information from real-world data and enables reverse color prediction from depth. Together with the proposed Syn-Real CutMix strategy, Back2Color effectively harmonizes joint training on synthetic and real datasets, improving the robustness of unsupervised monocular depth estimation in real traffic scenes.
  • We introduce Auto-learning Uncertainty Temporal-Spatial Fusion (Auto-UTSF), which robustly handles non-rigid motion and occlusions by adaptively fusing complementary spatial and temporal supervision through uncertainty estimation.
  • Extensive experiments on challenging urban driving benchmarks demonstrate that the proposed method consistently improves depth estimation accuracy while maintaining computational efficiency, and produces depth maps with clearer object boundaries.
@@
Figure 3: Color transformation for specific datasets: we showcase the capability of our Back2Color framework to effectively transform various synthetic datasets, including vKITTI2[4], SHIFT[51], and UrbanSyn[gómez2023one], into Syn2Real images. These images closely emulate the real-world styles of KITTI[10, 12] and Cityscapes[7], demonstrating high fidelity and credibility.
@@ II-A Unsupervised Depth Estimation
Unsupervised depth estimation aims to predict distances from stereo views or video sequences. Initial methods like those by Godard et al.[13] introduced disparity consistency loss, while Zhou et al.[67] pioneered depth and pose estimation simultaneously. Recent advancements[18, 57, 43, 66, 22, 38] refine network structures for higher accuracy. Techniques like PackNet[18] and CADepth[57] enhance geometric and appearance information propagation. HR-Depth[43] and DIFFNet[66] utilize HRNet for better resolution utilization. RA-Depth[22] employs HRNet for both encoding and decoding. In natural language processing, the Transformer architecture[55] has been successful, later adapted to computer vision tasks[8, 41]. It serves as the backbone for depth estimation[46, 54, 63, 62, 34], but faces challenges like computational demands and real-time constraints. Existing methods prioritize model architecture, often overlooking crucial obstacles and targets. Additionally, depth estimation models also face the issue of overfitting on specific datasets and a lack of continuous learning capability[60].
@@ II-C Style transfer in urban scenes
Research on style transfer for urban scene images has long been ongoing due to its importance in autonomous driving technology. Isola et al.[29] proposed a conditional GAN framework for generating urban scene images from semantic segmentation images of Cityscapes in 2017. However, generative models may produce random and plausible objects, disrupting the scene structure, as shown in their experiments. Atapour et al.[1] utilized style transfer and adversarial training to convert real-world color images into synthetic environment-style images. Huang et al.[25] applied GAN-based domain transfer for data augmentation in urban scene weather-time conversion. To minimize the gap between source and target domains, Gomez et al.[16] aligned colors across different domains in the Lab color space. Zheng et al.[65] used GANs to convert synthetic colors to real-world colors. With advancements in stable diffusion[61, 53], ControlNet [61] is now capable of generating images based on various conditions, including text, pose, Canny edges, and depth. This capability theoretically enables the generation of RGB images from urban depth maps, as demonstrated by RISHI et al. [53]. However, experiments [53] indicate that diffusion models do not strictly adhere to depth information and may arbitrarily alter target classes. Additionally, diffusion models require 2-5 seconds to generate a single image espically trainning with CNN model, and each condition demands extensive data for training. This makes it challenging to quickly fine-tune a depth estimation model for specific real-world scenarios. In contrast, our proposed ColorNet addresses these limitations, allowing for rapid deployment and real-time operation.
@@ III Methodology
In this section, we present the Back2Color framework, which captures the depth-to-color mapping within the target domain and transfers synthetic colors to pseudo-realistic ones (III-A1). Back2Color is further enhanced with Real-Syn CutMix (III-A2) to jointly train for unsupervised depth estimation in the real world and supervised depth estimation in synthetic data. Then, we describe the Auto-learning Uncertainty Temporal-Spatial Fusion (Auto-UTSF) module (III-B), designed to address challenges associated with non-rigid motion. Throughout this paper, Depth and Disparity are considered equivalent.
@@ III-A1 Transform depth to specific colors
Objectively speaking, predicting color from 3D structural information constitutes a one-to-many mapping, a non-standard conditional generative tasks, where the same 3D structure can correspond to a multitude of different color images. However, under specific conditions, this seemingly impossible task becomes feasible: Assuming that the color images in the target domain exhibit relatively high correlation. The colors in real word datasets like Cityscapes and KITTI are notably different, as shown in Fig.3. Therefore, in order to bridge the substantial gap between synthetic $\mathcal{S}\in$ {vKITTI2(vK)[4], SHIFT(SH)[51], UrbanSyn(Ur)[gómez2023one]} and real-world $\mathcal{X}\in$ {KITTI(KT)[10, 12], Cityscapes(Cs)[7]} datasets, we can learn the depth-to-color mapping in the target domain (real world color) and apply it to the source domain (synthetic color), thereby converting the source domain color to match those in the target domain. To bridge the appearance gap between synthetic and real-world data, we train a depth-to-color network $\mathcal{F}_{d2c}$ that predicts color images from estimated depth maps in the real-world domain. Given a predicted depth map $\hat{\mathbf{D}}_{\mathcal{X}}$ produced by the depth estimation network, the depth-to-color network reconstructs the corresponding color image as
$\hat{\mathbf{I}}_{\mathcal{X}}=\mathcal{F}_{d2c}(\hat{\mathbf{D}}_{\mathcal{X}};\boldsymbol{\theta}_{d2c}),$ (1)
where $\boldsymbol{\theta}_{d2c}$ denotes the learnable parameters of the depth-to-color network.
The reconstructed color image $\hat{\mathbf{I}}_{\mathcal{X}}$ is supervised by the corresponding real color image $\mathbf{I}_{\mathcal{X}}$ using a photometric reconstruction loss, defined as
$\mathcal{L}_{d2c}=pe(\hat{\mathbf{I}}_{\mathcal{X}},\mathbf{I}_{\mathcal{X}}),$ (2)
where $pe(\cdot,\cdot)$ denotes the photometric error function.
Next, we investigate whether the depth-to-color mapping learned from the real-world domain can be transferred to synthetic data. Specifically, after training the depth-to-color network $\mathcal{F}_{d2c}$ on real-world samples, we apply the learned model to predicted depth maps from the synthetic domain in order to generate color images that resemble the appearance characteristics of the target domain (e.g., KITTI), as illustrated in Fig. 3. Formally, given a predicted synthetic depth map $\hat{\mathbf{D}}_{\mathcal{S}}$, the corresponding color image is generated as
$\hat{\mathbf{I}}_{\mathcal{S}\rightarrow\mathcal{X}}=\mathcal{F}_{d2c}(\hat{\mathbf{D}}_{\mathcal{S}};\boldsymbol{\theta}_{d2c}),$ (3)
where $\hat{\mathbf{I}}_{\mathcal{S}\rightarrow\mathcal{X}}$ denotes the synthetic color image transformed to match the appearance of the real-world domain.
@@ III-A2 Syn-Real CutMix
The objective of CutMix in our framework is to construct enhanced training pairs $(\tilde{\mathbf{I}},\mathbf{D})$ for improving robustness across domains. Unlike the original CutMix [59], which mixes two independent image–label pairs, the image pair used in our setting, $(\mathbf{I}_{\mathcal{S}},\hat{\mathbf{I}}_{\mathcal{S}\rightarrow\mathcal{X}})$, represents two different color appearances corresponding to the same synthetic depth label $\mathbf{D}_{\mathcal{S}}$. Therefore, we perform mixing only in the color space while keeping the depth supervision unchanged.
To introduce spatial diversity, as illustrated in Fig. 4, we denote the image pair $(\mathbf{I}_{\mathcal{S}},\hat{\mathbf{I}}_{\mathcal{S}\rightarrow\mathcal{X}})$ as $(\mathbf{I}_{1},\mathbf{I}_{2})$, which are randomly swapped with equal probability. A rectangular region with height and width larger than half of the original image is then randomly cropped from $\mathbf{I}_{1}$ and replaced with the corresponding region from $\mathbf{I}_{2}$, yielding the mixed image
$\tilde{\mathbf{I}}=\mathbf{M}\odot\mathbf{I}_{1}+(1-\mathbf{M})\odot\mathbf{I}_{2},$ (4)
where $\mathbf{M}\in\{0,1\}^{H\times W}$ denotes a binary mask indicating the preserved and replaced regions, and $\odot$ represents element-wise multiplication.
The Syn-Real enhanced images $\tilde{\mathbf{I}}$ are then combined with real-world target images $\mathbf{I}_{\mathcal{X}}$ within the same training batch. This joint batching strategy allows the monocular depth estimation network to be trained simultaneously on real-world, synthetic, and Syn-to-Real samples. The spatial randomization introduced by Syn-Real CutMix further encourages the network to focus on structural cues rather than specific color appearances, thereby improving generalization across domains.
@@
Figure 6: Illustrates the complementary characteristics of temporal and spatial reprojection, which motivates the proposed Auto-UTSF strategy.
III-B Auto-learning Uncertainty Temporal–Spatial Fusion
During the evaluation of the Back2Color framework, we observe that the widely adopted auto-masking strategy [14] often fails in complex urban scenes, particularly on the Cityscapes dataset, as illustrated in Fig. 5. By inspecting the auto-masking outputs, we find that non-rigid moving objects, such as vehicles and pedestrians, cannot be reliably filtered. This limitation is further aggravated by the complexity of real-world traffic scenes and the strong fitting capability of modern depth networks, which may compensate for non-rigid motion by adjusting depth predictions to remain consistent with camera poses [27]. Moreover, masking strategies based on optical flow [69] completely fail for objects moving in the opposite direction of the ego-motion, as shown in Fig. 5.
In temporal sequences, freely moving objects violate the rigid-scene assumption underlying depth- and pose-based reprojection, as depicted in Fig. 6. Auto-masking attempts to identify such pixels by comparing reprojection errors before and after geometric warping:
$\mu_{i}=\mathbb{I}\!\left[\min_{k}pe(\mathbf{I}_{t},\mathbf{I}_{k\rightarrow t})<\min_{k}pe(\mathbf{I}_{t},\mathbf{I}_{k})\right],$ (5)
where $k\in\{t-1,t+1\}$. However, this heuristic is insufficient when depth predictions can be adjusted to rationalize non-rigid motion, leading to incorrect supervision.
Notably, non-rigid motion only affects temporally adjacent frames, whereas in spatially related stereo pairs captured at the same time, such motion does not occur. Therefore, instead of simply excluding unreliable pixels, introducing spatial (stereo) supervision provides a more direct and effective solution. Nevertheless, temporal and spatial reprojections exhibit complementary characteristics. Temporal reprojection benefits from larger baselines and multi-frame context, making it suitable for distant regions and occlusion completion, while spatial reprojection offers reliable supervision for non-rigid objects but suffers from limited baseline and asymmetric occlusions. These observations motivate a unified strategy that adaptively fuses temporal and spatial supervision.